# Enhancing Content-based Recommendation via Large Language Model

### v1

@@ Abstract.
In real-world applications, users express different behaviors when they interact with different items, including implicit click/like interactions, and explicit comments/reviews interactions. Nevertheless, almost all recommender works are focused on how to describe user preferences by the implicit click/like interactions, to find the synergy of people. For the content-based explicit comments/reviews interactions, some works attempt to utilize them to mine the semantic knowledge to enhance recommender models. However, they still neglect the following two points: (1) The content semantic is a universal world knowledge; how do we extract the multi-aspect semantic information to empower different domains? (2) The user/item ID feature is a fundamental element for recommender models; how do we align the ID and content semantic feature space? In this paper, we propose a ‘plugin’ semantic knowledge transferring method LoID, which includes two major components: (1) LoRA-based large language model pretraining to extract multi-aspect semantic information; (2) ID-based contrastive objective to align their feature spaces. We conduct extensive experiments with SOTA baselines on real-world datasets, the detailed results demonstrating significant improvements of our method LoID.
@@
ccs: Information systems Recommender systemsccs: Computing methodologies Neural networks
@@ 1. Introduction
Background. With the boom of digital information, billions of user requests are produced daily, so recommender systems (RSs) have become an integral part of Internet platforms. To capture users’ interests more accurately, RSs have gone through several milestones, such as the logistic regression with hand-crafted features (e.g., FM (Rendle, 2010)), the neural networks (e.g., WideDeep (Cheng et al., 2016), YoutubeNet (Covington et al., 2016)), the sequential signal (e.g., DIN (Zhou et al., 2018), SIM (Pi et al., 2020)), and the multi-hop graph signal (e.g., PinSage (Ying et al., 2018), DGRec (Liangwei Yang, 2023)). In retrospect, these effective methods are based on the collaborative filtering (CF) idea and extend the boundaries of RSs. However, the CF framework also limits them under the case of cold-start and data sparsity problems. The reason is that the CF idea aims to mine the user/item pattern intelligence from data, to discover and recommend high-click candidate items for a user while it is hard to understand the users’ fine-grained and multi-aspect interests. In fact, instead of mining user preferences from massive user-item interaction logs, the user always leaves some reviews/comments to explain further his/her feelings about this interaction, which provides an explicit way to understand the users’ complex interests in language semantic space.
@@
    Pretraining semantic ‘plugin’: On the one hand, the ideal semantic information should act as a universal role to support all domains. On the other hand, different domains have their main aspects of semantic information that are related to recommendation. Considering such a trade-off, we borrow the LoRA strategy idea to train a small set of parameters for each domain, which could serve as plugins that can be added seamlessly to the target domain without further re-training.
@@
  • We give a ‘plugin’ idea to transfer semantic knowledge, which sheds light on building a new paradigm for recommendation.
@@ 2.1. Problem Statement
This work considers a brief task: For one domain dataset $\mathcal{D}$, it contains user-item ratings and textual contents, and each user-item interaction primarily comprises the following four elements: user ${\mathrm{u}}$, item ${\mathrm{i}}$, the corresponding rating $r_{\mathrm{u,i}}$, and the textual content token list $\textbf{t}_{\mathrm{u,i}}$ left by the user. Our model aims to predict ratings by historical content and user/item ID information.
Besides, to test the ‘plugin’ idea effectiveness, we further consider multi-domain scenarios, that is: given source domain content information $\{(\textbf{t}^{\mathrm{source_{1}}},r^{\mathrm{source_{1}}}),(\textbf{t}^{\mathrm{source_{2}}},r^{\mathrm{source_{2}}}),\dots\}$, predict target domain $\mathcal{D}_{\mathrm{target}}$ rating score.
@@ 2.2. Low-Rank Adaptation (LoRA)
Before going on, we first explain the LoRA (Hu et al., 2021), to show the basic idea of how to fast tuning an LLM (Devlin et al., 2018). Indeed, the unit block of LLM, transformer (Vaswani et al., 2017), consists of two parts: (Masked) attention and MLPs (FFM). The two parts introduce four pre-trained parameter matrices (e.g., $\mathbf{W}_{\mathrm{Q}},\mathbf{W}_{\mathrm{K}},\mathbf{W}_{\mathrm{V}},\mathbf{W}_{\mathrm{FFM}}$). Then, the challenge is: how to update those matrices without re-training them or involving new large parameters. Thereby, LoRA was proposed by assigning two small matrics for any pre-trained matrix $\mathbf{W}_{\mathrm{BERT}}^{\mathrm{Frozen}}\in\theta^{\mathrm{Frozen}}_{\mathrm{BERT}}$:
(1) $\small\begin{split}\mathbf{W}^{\mathrm{LoRA}}=\mathbf{W}_{\mathrm{BERT}}^{\mathrm{Frozen}}+\mathbf{B}^{\mathrm{LoRA}}\mathbf{A}^{\mathrm{LoRA}},\end{split}$
@@
(2) $\small\begin{split}\theta^{\mathrm{LoRA}}_{\mathrm{Task}}=\mathrm{LoRA(}\theta^{\mathrm{Frozen}}_{\mathrm{BERT}},\mathrm{Task}),\quad\theta_{\mathrm{Task}}=\theta^{\mathrm{Frozen}}_{\mathrm{BERT}}+\theta^{\mathrm{LoRA}}_{\mathrm{Task}},\end{split}$
where $\theta^{\mathrm{LoRA}}_{\mathrm{Task}}$ is extra added parameters and $\theta_{\mathrm{Task}}$ is all parameters [^1] In our work, we only introduce extra 4M $\theta^{\mathrm{LoRA}}_{\mathrm{Task}}$ parameters to tune 110M $\theta^{\mathrm{pre}}_{\mathrm{BERT}}$ BERT..
@@ 3.1. Overview
The architecture of our method LoID is illustrated in Figure 1, which includes two processes. In part (a), we first train the LoRA parameters of source domain as ”plug-ins” to enhance target domain prediction without further re-training. In part (b), we first extract historical contents of user/item to obtain user/item semantic representation, and then align the ID and semantic to make target domain rating prediction, note that the source domain LoRA plugin is an optional choice.
3.2. LoRA-based Encoder Pretraining
@@
Source LoRA. To get users’ behavior data from the source domain, We leverage the LoRA strategy to pre-train the source domain [^2] For better understanding, we use BERT as LLM to show our method details.. In the pre-training task, we leverage rating prediction by $\mathbf{CLS}_{\mathrm{source}}$ token to predict rating directly:
@@
(3) $\small\begin{split}\theta^{\mathrm{LoRA}}_{\mathrm{source}}=\mathrm{LoRA(}\theta^{\mathrm{Frozen}}_{\mathrm{BERT}},\mathrm{Task}_{\mathrm{source}})&,\ \ \ \theta_{\mathrm{source}}=\theta^{\mathrm{Frozen}}_{\mathrm{BERT}}+\theta^{\mathrm{LoRA}}_{\mathrm{source}},\\ \mathbf{CLS}^{\mathrm{source}}_{\mathrm{u,i}}=\text{BERT}(\textbf{t}_{\mathrm{u,i}}^{\mathrm{source}},\theta_{\mathrm{source}}),&\quad\hat{r}_{\mathrm{u,i}}^{\mathrm{source}}=\text{Predict}(\mathbf{CLS}^{\mathrm{source}}_{\mathrm{u,i}}),\\ \end{split}$
@@
where BERT()($\cdot$)( ⋅ ) is the LLM forward procedure, Predict()($\cdot$)( ⋅ ) is an MLP to generate the prediction score $\hat{r}_{\mathrm{u,i}}^{\mathrm{source}}$, then we adopt the Mean Squared Error (MSE) loss to optimize LoRA parameters:
@@
(4) $\small\begin{split}\mathcal{L}_{\mathrm{pre-train}}^{\mathrm{source}}=\frac{1}{|\mathcal{D}_{\mathrm{source}}|}\sum_{\mathrm{(u,i)}}^{\mathcal{D}_{\mathrm{source}}}(\hat{r}_{\mathrm{u,i}}^{\mathrm{source}}-r_{\mathrm{u,i}}^{\mathrm{source}})^{2},\end{split}$
@@
where $|\mathcal{D}_{\mathrm{source}}|$ means the number of the samples of source domains, $\hat{r}_{\mathrm{u,i}}$/$r_{\mathrm{u,i}}$ represent predicted/real rating respectively. After pre-training all source domain’s LoRA parameters, we can use them as plugins to enhance BERT when dealing with other target domain tasks without further re-training.
3.3. Re-LoRA In Target Domain
Target LoRA. Enabling the model to adapt to the target domain, we introduce a LoRA module (i.e., Target LoRA) again in the target domain to fine-tune the target LoRA parameters. In the Re-LoRA process, we freeze the parameters of the BERT and source LoRA, and the only part that needs training is Target LoRA:
@@ 3.3. Re-LoRA In Target Domain
(5) $\footnotesize\begin{split}\mathbf{W}^{\mathrm{LoRA}}_{\mathrm{target}}&=\mathbf{W}_{\mathrm{BERT}}^{\mathrm{Frozen}}+\mathbf{B}_{\mathrm{source}}^{\mathrm{LoRA,Frozen}}\mathbf{A}_{\mathrm{source}}^{\mathrm{LoRA,Frozen}}+\mathbf{B}_{\mathrm{target}}^{\mathrm{LoRA}}\mathbf{A}_{\mathrm{target}}^{\mathrm{LoRA}},\\ \theta_{\mathrm{target}}=\{\theta^{\mathrm{Frozen}}_{\mathrm{BERT}}&,\theta^{\mathrm{LoRA,Frozen}}_{\mathrm{source}}\}+\mathrm{LoRA(}\{\theta^{\mathrm{Frozen}}_{\mathrm{BERT}},\theta^{\mathrm{LoRA,Frozen}}_{\mathrm{source}}\},\mathrm{Task}_{\mathrm{target}}),\end{split}$
where $\theta_{\mathrm{target}}$ is the LLM part’s parameters of target domain. In the next, we explain how to align the language semantic and ID space.
@@ 3.4. ID-based Contrastive Learning
As the fundamental element of CF-based recommender work, the user/item IDs are indispensable in achieving personalized signals. How to align the space of semantic/ID is the key to making the content information more competitive for industrial RS.
@@
(6) $\small\begin{split}\mathbf{CLS}_{\mathrm{u}}^{\mathrm{k}}=\text{BERT}(\textbf{t}_{\mathrm{u}}^{\mathrm{k}},\theta_{\mathrm{target}}),\quad\mathbf{CLS}_{\mathrm{i}}^{\mathrm{k}}=\text{BERT}(\textbf{t}_{\mathrm{i}}^{\mathrm{k}},\theta_{\mathrm{target}}),\end{split}$
where $\mathbf{CLS}_{\mathrm{u}}^{\mathrm{k}}$ and $\mathbf{CLS}_{\mathrm{i}}^{\mathrm{k}}$ are the embeddings for the user’s and item’s $k^{th}$ content.$\textbf{t}_{\mathrm{u}}^{\mathrm{k}}$ and $\textbf{t}_{\mathrm{i}}^{\mathrm{k}}$ are the corresponding textual content token list, and $\theta_{\mathrm{target}}$ stands for the LLM parameter on the target domain.
Attention Layer. On top of the user/item historical content $\mathbf{CLS}$ semantic information, we devise a novel attention mechanism to exchange the user/item semantic/ID information. Specifically, we integrate the semantic information $\{\mathbf{CLS}_{\mathrm{u}}^{1},\dots,\mathbf{CLS}_{\mathrm{u}}^{k}\}$ into the item representation $\mathbf{p}_{i}$, versus versa for user representation $\mathbf{p}_{u}$. By this mechanism, ID embedding and $\mathbf{CLS}$ token are fused, after which we obtain the updated item representation $\mathbf{v}_{i}$. Similarly, on the user side, we obtain ${\mathbf{v}_{u}}$ as follows:
(7) $\small\begin{split}\mathbf{v}_{i}&=\text{Attention}(\mathbf{p}_{i},\{\mathbf{CLS}_{\mathrm{u}}^{\mathrm{1}},\dots,\mathbf{CLS}_{\mathrm{u}}^{\mathrm{k}}\},\{\mathbf{CLS}_{\mathrm{u}}^{\mathrm{1}},\dots,\mathbf{CLS}_{\mathrm{u}}^{\mathrm{k}}\}),\\ \mathbf{v}_{u}&=\text{Attention}(\mathbf{p}_{u},\{\mathbf{CLS}_{\mathrm{i}}^{\mathrm{1}},\dots,\mathbf{CLS}_{\mathrm{i}}^{\mathrm{k}}\},\{\mathbf{CLS}_{\mathrm{i}}^{\mathrm{1}},\dots,\mathbf{CLS}_{\mathrm{i}}^{\mathrm{k}}\}),\\ \end{split}$
@@
Contrastive Loss. The Attention mechanism primarily focuses on different parts of the input sequence but does not explicitly consider the relationship between users and items. So, we conduct contrastive learning, enhancing the similarity between interactive users and items. For each updated representation $\mathbf{v}_{i}$/$\mathbf{v}_{u}$, treated as an anchor, we pair it with the original user/item representation as a positive sample $\mathbf{p}^{+}_{u}/\mathbf{p}^{+}_{i}$, and another representation in the batch as a negative sample $\mathbf{p}^{-}_{u}/\mathbf{p}^{-}_{i}$. The goal is to minimize the distance $\mathrm{D}(anchor,positive)$ and maximize $\mathrm{D}(anchor,negative)$. The loss function is minimized to ensure that the representations of similar items/users are closer in terms of the distance $\mathrm{D}$.
(8) $\small\begin{split}\mathrm{D}&=||anchor-sample||^{2}_{2},\\ \mathcal{L}_{\mathrm{cl}}=\max(&0,\Delta+\mathrm{D}(\mathbf{v}_{i},\mathbf{p}^{+}_{u})-\mathrm{D}(\mathbf{v}_{i},\mathbf{p}^{-}_{u}))\\ +\max(&0,\Delta+\mathrm{D}(\mathbf{v}_{u},\mathbf{p}^{+}_{i})-\mathrm{D}(\mathbf{v}_{u},\mathbf{p}^{-}_{i})),\\ \end{split}$
where ${\mathbf{v}_{i}}/\mathbf{v}_{u}$ is the representation of the updated item/user, ${\mathbf{p}^{+}_{u}/\mathbf{p}^{+}_{i}}$ represents the original feature representation of the user/item associated with the current content, as positive sample. ${\mathbf{p}^{-}_{u}/\mathbf{p}^{-}_{i}}$ represents the negative sample. ${\Delta}$ is a margin constant that enforces a ”safe distance” constraint on correctly classified samples, ensuring that the model’s classification results have sufficient robustness.

### v2

@@ Abstract.
In real-world applications, users express different behaviors when they interact with different items, including implicit click/like interactions, and explicit comments/reviews interactions. Nevertheless, almost all recommender works are focused on how to describe user preferences by the implicit click/like interactions, to find the synergy of people. For the content-based explicit comments/reviews interactions, some works attempt to utilize them to mine the semantic knowledge to enhance recommender models. However, they still neglect the following two points: (1) The content semantic is a universal world knowledge; how do we extract the multi-aspect semantic information to empower different domains? (2) The user/item ID feature is a fundamental element for recommender models; how do we align the ID and content semantic feature space? In this paper, we propose a ‘plugin’ semantic knowledge transferring method LoID, which includes two major components: (1) LoRA-based large language model pretraining to extract multi-aspect semantic information; (2) ID-based contrastive objective to align their feature spaces. We conduct extensive experiments with SOTA baselines to demonstrate superiority of our method LoID.
@@
§The publication does not relate to the author’s work at Kuaishou.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679913isbn: 979-8-4007-0436-9/24/10ccs: Information systems Recommender systems
@@ 1. Introduction
Background. With the boom of digital information, billions of user requests are produced daily, so recommender systems (RSs) have become an integral part of Internet platforms. To capture users’ interests more accurately, RSs have gone through several milestones, such as the logistic regression with hand-crafted features (e.g., FM (Rendle, 2010)), the neural networks (e.g., WideDeep (Cheng et al., 2016), YoutubeNet (Covington et al., 2016)), the sequential signal (e.g., DIN (Zhou et al., 2018), SIM (Pi et al., 2020)), and the multi-hop graph signal (e.g., PinSage (Ying et al., 2018), DGRec (Liangwei Yang, 2023)). In retrospect, these effective methods are based on the collaborative filtering (CF) idea and extend the boundaries of RSs. However, the CF framework also limits them under the case of cold-start and data sparsity problems. The reason is that the CF idea aims to mine the user/item pattern intelligence from data, to discover and recommend high-click candidate items for a user while it is hard to understand the users’ fine-grained and multi-aspect interests. In fact, instead of mining user preferences from massive user-item interaction logs, the user always leaves some reviews/comments to explain further his/her feelings about this interaction, which provides an explicit way to understand the users’ complex interests in language semantic space. And the recent effort MoRec (Yuan et al., 2023) claimed that multi-modal information yield better results in sequential recommendation.
@@
    Pretraining semantic ‘plugin’: On the one hand, the ideal semantic information should act as a universal role to support all domains. On the other hand, different domains have their main aspects of semantic information that are related to recommendation. We borrow the LoRA strategy idea to train a small set of parameters for each domain, which could serve as plugins. Then we fused them by DARE  (Yu et al., 2023)without further re-training.
@@
  • We give a ‘plugin’ idea to transfer semantic knowledge, which sheds light on building a new paradigm for recommendation [^1] Codes at https://github.com/cjx96/LoID..
@@ 2.1. Problem Statement
This work considers a brief task: For one domain dataset $\mathcal{D}$, and each user-item interaction primarily comprises the following four elements: user ${\mathrm{u}}$, item ${\mathrm{i}}$, the corresponding rating $r_{\mathrm{u,i}}$, and the textual content token list $\textbf{t}_{\mathrm{u,i}}$ left by the user. Our model aims to predict ratings by historical content and user/item ID information.
Besides, to test the effectiveness of the ‘plugin’ idea , we further consider multi-domain scenarios, that is: given source domain content information $\{(\textbf{t}^{\mathrm{source_{1}}},r^{\mathrm{source_{1}}}),(\textbf{t}^{\mathrm{source_{2}}},r^{\mathrm{source_{2}}}),\dots\}$, predict target domain $\mathcal{D}_{\mathrm{target}}$ rating score.
@@ 2.2. Low-Rank Adaptation (LoRA)
Before going on, we first explain the LoRA (Hu et al., 2021), to show the basic idea of how to fast tuning an LLM (Devlin et al., 2018). Indeed, the unit block of LLM, transformer (Vaswani et al., 2017), consists of two parts: (Masked) attention and MLPs (FFN). The two parts introduce four pre-trained parameter matrices (e.g., $\mathbf{W}_{\mathrm{Q}},\mathbf{W}_{\mathrm{K}},\mathbf{W}_{\mathrm{V}},\mathbf{W}_{\mathrm{FFN}}$). Then, the challenge is: how to update those matrices without re-training them or involving new large parameters. Thereby, LoRA was proposed by assigning two small matrices for any pre-trained matrix $\mathbf{W}_{\mathrm{LLM}}^{\mathrm{Frozen}}\in\theta^{\mathrm{Frozen}}_{\mathrm{LLM}}$:
(1) $\footnotesize\begin{split}\mathbf{W}^{\mathrm{LoRA}}=\mathbf{W}_{\mathrm{LLM}}^{\mathrm{Frozen}}+\mathbf{B}^{\mathrm{LoRA}}\mathbf{A}^{\mathrm{LoRA}},\end{split}$
@@
(2) $\footnotesize\begin{split}\theta^{\mathrm{LoRA}}_{\mathrm{Task}}=\mathrm{LoRA(}\theta^{\mathrm{Frozen}}_{\mathrm{LLM}},\mathrm{Task}),\quad\theta_{\mathrm{Task}}=\theta^{\mathrm{Frozen}}_{\mathrm{LLM}}+\theta^{\mathrm{LoRA}}_{\mathrm{Task}},\end{split}$
where $\theta^{\mathrm{LoRA}}_{\mathrm{Task}}$ is extra added parameters and $\theta_{\mathrm{Task}}$ is all parameters [^2] In our work, we only introduce extra 4M $\theta^{\mathrm{LoRA}}_{\mathrm{Task}}$ parameters to tune 110M $\theta^{\mathrm{pre}}_{\mathrm{BERT}}$ BERT.. Among different domains, $\theta^{\mathrm{LoRA}}_{\mathrm{Task}}$ has the same structure, so it is a proper semantic component for knowledge transfer  (Fu et al., 2024).
2.3. Drop and Rescale (DARE)
Assuming that we have fine-tuned the same LLM in different specific tasks and obtained several (supervised fine-tuned) SFT models as $\{\theta_{\mathrm{Task_{1}}},\theta_{\mathrm{Task_{2}}},\dots\}$, a challenge is: how to merge them as an omnipotent LLM? To answer it, DARE (Yu et al., 2023) was proposed, which devises a drop-then-rescale strategy on the gradient update sets $\{\delta_{\mathrm{Task_{1}}},\delta_{\mathrm{Task_{2}}},\cdots\}$ to make the multiple homogenous LLM merging, here we take the LoRA-Style SFT as an example:
(3) $\footnotesize\begin{split}\delta_{\mathrm{Task_{i}}}=\theta_{\mathrm{Task_{i}}}-\theta_{\mathrm{LLM}}=\mathbf{B}^{\mathrm{LoRA}}_{\mathrm{Task_{i}}}\mathbf{A}^{\mathrm{LoRA}}_{\mathrm{Task_{i}}}\\ \end{split}$
Next, DARE performs random drop on $\delta_{\mathrm{Task}}$ with probability $p$, and then rescales the remaining ones by a factor $1/(1-p)$.
(4) $\footnotesize\begin{split}\tilde{\delta}_{\mathrm{Task_{i}}}=\mathrm{Drop}(\delta_{\mathrm{Task_{i}}}),\quad\tilde{\delta}^{\mathrm{DARE}}_{\mathrm{Task_{i}}}=\tilde{\delta}_{\mathrm{Task_{i}}}/(1-p)\end{split}$
where $p$ can reach 0.9 even 0.99, $\tilde{\delta}_{\mathrm{Task_{i}}}$ and $\tilde{\delta}^{\mathrm{DARE}}_{\mathrm{Task_{i}}}$ represent the delta parameters after dropping and rescaling respectively.
(5) $\footnotesize\begin{split}\theta^{\mathrm{DARE}}_{\mathrm{LLM}}&=\mathrm{DARE(}\theta^{\mathrm{Frozen}}_{\mathrm{LLM}},\theta_{\mathrm{Task_{1}}},\theta_{\mathrm{Task_{2}}},\cdots,\theta_{\mathrm{Task_{n}}})\\ &=\theta_{\mathrm{LLM}}^{\mathrm{Frozen}}+\tilde{\delta}^{\mathrm{DARE}}_{\mathrm{Task_{1}}}+\tilde{\delta}^{\mathrm{DARE}}_{\mathrm{Task_{2}}}+\tilde{\delta}^{\mathrm{DARE}}_{\mathrm{Task_{3}}}+\cdots\end{split}$
where $\theta^{\mathrm{DARE}}_{\mathrm{LLM}}$ is the merged SFT LLM.
@@ 3.1. Overview
The architecture of our method LoID is illustrated in Figure 1. In part (a), we train the LoRA parameters of source domains as ”plug-ins” to enhance target domain prediction without further re-training. In part (b), we extract historical contents of user/item to obtain user/item semantic representation, and then align the ID and semantic to make target domain rating prediction, note that the source domain LoRA plugin is an optional choice.
3.2. Source Domain LoRA Pretraining
Figure 1. Model Architecture of the proposed LoID.
@@ 3.2. Source Domain LoRA Pretraining
Source LoRA. To get users’ behavior data from the source domain, We leverage the LoRA strategy to pre-train the source domain. In the pre-training task, we leverage rating prediction by $\mathbf{CLS}_{\mathrm{source}}$ token to predict rating directly, here we take BERT as default LLM:
@@
$\footnotesize\begin{split}\theta^{\mathrm{LoRA}}_{\mathrm{source_{k}}}=\mathrm{LoRA(}\theta^{\mathrm{Frozen}}_{\mathrm{BERT}},\mathrm{Task}_{\mathrm{source_{k}}})&,\ \ \ \theta_{\mathrm{source_{k}}}=\theta^{\mathrm{Frozen}}_{\mathrm{BERT}}+\theta^{\mathrm{LoRA}}_{\mathrm{source_{k}}},\\ \mathbf{CLS}^{\mathrm{source}}_{\mathrm{u,i}}=\text{BERT}(\textbf{t}_{\mathrm{u,i}}^{\mathrm{source_{k}}},\theta_{\mathrm{source_{k}}}),&\quad\hat{r}_{\mathrm{u,i}}^{\mathrm{source_{k}}}=\text{Predict}(\mathbf{CLS}^{\mathrm{source_{k}}}_{\mathrm{u,i}}),\\ \end{split}$
@@
where $k$ denotes the $k^{th}$ domain in the source domains. where BERT()($\cdot$)( ⋅ ) is the LLM forward procedure, Predict()($\cdot$)( ⋅ ) is an MLP to generate the prediction score $\hat{r}_{\mathrm{u,i}}^{\mathrm{source}}$, then we adopt the Mean Squared Error (MSE) loss to optimize LoRA parameters:
@@
(6) $\footnotesize\begin{split}\mathcal{L}_{\mathrm{pre-train}}^{\mathrm{source}}=\frac{1}{|\mathcal{D}_{\mathrm{source}}|}\sum_{\mathrm{(u,i)}}^{\mathcal{D}_{\mathrm{source}}}(\hat{r}_{\mathrm{u,i}}^{\mathrm{source}}-r_{\mathrm{u,i}}^{\mathrm{source}})^{2},\end{split}$
@@
where $|\mathcal{D}_{\mathrm{source}}|$ means the number of the samples of source domains, $\hat{r}_{\mathrm{u,i}}$/$r_{\mathrm{u,i}}$ represent predicted/real rating respectively.
3.3. Target Domain Re-LoRA
After obtaining several source semantic ‘plugin’ LoRAs, we then transfer their general modality knowledge to target in two steps.
Source LoRA Merging. The first step is to merge these pre-trained source plugins with LLM, as discussed before, we follow DARE (Yu et al., 2023) idea to merge multiple LoRAs parameters.
@@ 3.3. Target Domain Re-LoRA
(7) $\footnotesize\begin{split}\theta^{\mathrm{DARE}}_{\mathrm{BERT}}=\mathrm{DARE(}\theta^{\mathrm{Frozen}}_{\mathrm{BERT}},\theta_{\mathrm{source_{1}}},\theta_{\mathrm{source_{2}}},\cdots,\theta_{\mathrm{source_{n}}})\end{split}$
where $n\geq 0$ represents the avaliable source domains, note that the $n=0$ means ignoring all source domain LoRAs.
Target LoRA. Enabling the model to adapt to the target domain, we introduce a LoRA module (i.e., Target LoRA) again in the target domain. In the Re-LoRA, only the Target LoRA tuning its parameters; the merged BERT parameters are frozen:
$\footnotesize\begin{split}\theta^{\mathrm{LoRA}}_{\mathrm{target}}=\mathrm{LoRA(}\theta^{\mathrm{DARE,Frozen}}_{\mathrm{BERT}},\mathrm{Task}_{\mathrm{target}})&,\ \ \ \theta_{\mathrm{target}}=\theta^{\mathrm{DARE,Frozen}}_{\mathrm{BERT}}+\theta^{\mathrm{LoRA}}_{\mathrm{target}},\\ \end{split}$
where $\theta_{\mathrm{target}}$ is the LLM part’s parameters of target domain.
@@ 3.4. ID-based Contrastive Learning
As the fundamental element of CF-based recommender work, the user/item IDs are indispensable in achieving personalized signals. How to align the independent feature space of semantic/ID is the key to making the content information more competitive for industrial RS. Therefore, we design an auxiliary task.
@@
(8) $\footnotesize\begin{split}\mathbf{CLS}_{\mathrm{u}}^{\mathrm{k}}=\text{BERT}(\textbf{t}_{\mathrm{u}}^{\mathrm{k}},\theta_{\mathrm{target}}),\quad\mathbf{CLS}_{\mathrm{i}}^{\mathrm{k}}=\text{BERT}(\textbf{t}_{\mathrm{i}}^{\mathrm{k}},\theta_{\mathrm{target}}),\end{split}$
where $\mathbf{CLS}_{\mathrm{u}}^{\mathrm{k}}$ and $\mathbf{CLS}_{\mathrm{i}}^{\mathrm{k}}$ are the embeddings for the user’s and item’s $k^{th}$ content.$\textbf{t}_{\mathrm{u}}^{\mathrm{k}}$ and $\textbf{t}_{\mathrm{i}}^{\mathrm{k}}$ are the corresponding textual contents.
Attention Layer. On top of the user/item historical content $\mathbf{CLS}$ semantic information, we devise a novel attention mechanism to exchange the user/item semantic/ID information. Specifically, we integrate the semantic information $\{\mathbf{CLS}_{\mathrm{u}}^{1},\dots,\mathbf{CLS}_{\mathrm{u}}^{k}\}$ into the item representation $\mathbf{p}_{i}$, versus versa for user representation $\mathbf{p}_{u}$.
(9) $\footnotesize\begin{split}\mathbf{v}_{i}&=\text{Attention}(\mathbf{p}_{i},\{\mathbf{CLS}_{\mathrm{u}}^{\mathrm{1}},\dots,\mathbf{CLS}_{\mathrm{u}}^{\mathrm{k}}\},\{\mathbf{CLS}_{\mathrm{u}}^{\mathrm{1}},\dots,\mathbf{CLS}_{\mathrm{u}}^{\mathrm{k}}\}),\\ \mathbf{v}_{u}&=\text{Attention}(\mathbf{p}_{u},\{\mathbf{CLS}_{\mathrm{i}}^{\mathrm{1}},\dots,\mathbf{CLS}_{\mathrm{i}}^{\mathrm{k}}\},\{\mathbf{CLS}_{\mathrm{i}}^{\mathrm{1}},\dots,\mathbf{CLS}_{\mathrm{i}}^{\mathrm{k}}\}),\\ \end{split}$
@@
Contrastive Loss. The Attention mechanism primarily focuses on different parts of the input sequence, but it does not consider the relationship between users and items. So, we conduct contrastive learning, enhancing the similarity between interactive users and items. For each updated representation $\mathbf{v}_{i}$/$\mathbf{v}_{u}$, treated as an anchor, we pair it with the original user/item feature representation as a positive sample $\mathbf{p}^{+}_{u}/\mathbf{p}^{+}_{i}$, and another representation in the batch as a negative sample $\mathbf{p}^{-}_{u}/\mathbf{p}^{-}_{i}$. The goal is to minimize the distance $\mathrm{D}(anchor,positive)$ and maximize $\mathrm{D}(anchor,negative)$.
$\footnotesize\begin{split}\mathrm{D}&=||anchor-sample||^{2}_{2},\\ \mathcal{L}_{\mathrm{cl}}=\max(0,\Delta+\mathrm{D}(\mathbf{v}_{i},\mathbf{p}^{+}_{u})&-\mathrm{D}(\mathbf{v}_{i},\mathbf{p}^{-}_{u}))+\max(0,\Delta+\mathrm{D}(\mathbf{v}_{u},\mathbf{p}^{+}_{i})-\mathrm{D}(\mathbf{v}_{u},\mathbf{p}^{-}_{i})),\\ \end{split}$
where ${\mathbf{v}_{i}}/\mathbf{v}_{u}$ is the representation of the updated item/user, ${\mathbf{p}^{+}_{u}/\mathbf{p}^{+}_{i}}$ represents the positive sample, and ${\mathbf{p}^{-}_{u}/\mathbf{p}^{-}_{i}}$ represents the negative sample. ${\Delta}$ is a margin constant that enforces a ”safe distance” constraint on correctly classified samples.